# The odd-parity altermagnetism: A spin group study

### 30 Mar 2026 v4

@@ Abstract
Following recent intensive studies on altermagnetism(ALM) characterized by non-relativistic even-parity spin splitting, realizing unconventional odd-parity magnetism has also attracted increasing interest. Here, using symmetry arguments based on spin-group analyses, we elucidate sufficient conditions for the emergence of odd-parity spin splitting in collinear antiferromagnetic systems, which is further established as the standard odd-parity ALM. It is derived that the odd-parity ALM arises from the following criteria: (i)the breaking nonmagnetic time reversal symmetry(TRS), i.e., the breaking real-space TRS; (ii)the long-range collinear compensated magnetism; (iii)the symmetry $C_{2}||\bar{E}]$ or its variant $C_{2}||\bar{E}\bm{\tau}]$ connecting opposite-spin sublattices, where $C_{2}$, $\bar{E}$, and $\bm{\tau}$ respectively represent a $180^{\circ}$ rotation around the axis perpendicular to spins, the inversion, and the translation between opposite-spin sublattices. Moreover, we utilize the well-known Haldane-Hubbard model to identify odd-parity spin splitting in the collinear ALM ground state, where (i)the nonmagnetic TRS is broken by opposite sublattice currents coming from the Haldane hopping; (ii)the symmetry $C_{2}||\bar{E}]$ is ensured because the currents flowing on opposite-spin sublattices are reversed.
@@
However, for collinear antiferromagnets with the breaking nonmagnetic TRS that can be caused by sublattice currents[38], light[33, 32], or the phase-dependent hopping coming from orbital orders[43], the symmetry $\bar{C}_{2}||T]$ appearing in even-parity ALM’s no longer holds. In contrast, the identity element in the nontrivial spin group $\bm{R}_{s}$ for collinear magnets, $\bar{C}_{2}||E]$, remains with $E$ being the identity in real space. Furthermore, as shown by Eq.(1), the odd-parity spin splitting in collinear magnets is ensured by the symmetry $C_{2}||C_{2z}]$ or its variant $C_{2}||C_{2z}\bm{\tau}]$, while for two-dimension systems, the operation $C_{2z}$ is equivalent to $\bar{E}$, such that the odd-parity spin splitting in two-dimension magnetic systems with long-range compensated collinear magnetism is ensured by the symmetry $C_{2}||\bar{E}]$ or $C_{2}||\bar{E}\bm{\tau}]$. Therefore, in strong contrast to even-parity ALM’s, the long-range compensated collinear magnets exhibiting odd-parity spin splitting have the following characteristics: (i)the breaking nonmagnetic TRS then the breaking inversion symmetry of the Bravais lattice; (ii)the symmetry $C_{2}||\bar{E}]$ or $C_{2}||\bar{E}\bm{\tau}]$, which directly enables that the nontrivial spin group expressed as $\bm{R}_{s}^{IV}=\{E||\bm{H}\}+\{C_{2}||A\bm{H}\}=\{E||\bm{H}\}+\{C_{2}||\bm{G}-\bm{H}\}$ has the distinct coset $A\bm{H}$ of the halved crystallographic Laue group $\bm{H}$ containing the inversion $\bar{E}$ or its variant such as $\bar{E}\bm{\tau}$. To sum up, in spit of the similar form of the nontrivial spin group to even-parity ALM’s, the odd-parity spin splitting in compensated collinear magnets need to be classified into a new type, i.e., odd-parity altermagnetism, not only because the symmetry $\bar{C}_{2}||T]$ is absent, but also the inversion symmetry $\bar{E}$ or its variant $\bar{E}\bm{\tau}$ is now contained in the coset of the halved subgroup $\bm{H}$, i.e., $A\bm{H}$.
@@
Haldane-Hubbard model.We use Haldane-Hubbard(HH) model[48] as an example where the nonmagnetic TRS is broken by sublattice currents to verify sufficient conditions for the occurrence of odd-parity ALM. As shown in Figs.2(a) and (b), the nonmagnetic TRS is broken by opposite currents flowing on two sublattices in the square- and hexagon-lattice magnetic system. Furthermore, because of the opposite current and compensated collinear magnetization on two sublattices, there emerges respectively a new form of symmetry $C_{2}||\bar{E}\bm{\tau}]$ and $C_{2}||\bar{E}]$ with $\bm{\tau}$ being the minimal vector connecting opposite-spin sublattices. Specifically, as shown in Fig.2(a), the symmetry $C_{2}||\bar{E}\bm{\tau}]$ connects opposite-spin sublattices with $\bm{\tau}=\hat{x}$, $\hat{y}$ being the unit vector along the $x$ and $y$ direction, respectively. However, in the hexagon-lattice magnetic system shown in Fig.2(b), two sublatttices are connected by the inversion $\bar{E}$ in the paramagnetic state, such that as the compensated collinear spin arrangement sets in, opposite-spin sublattices are associated by the symmetry $C_{2}||\bar{E}]$. Therefore, sufficient conditions for the appearance of odd-parity ALM have been reached for magnetic systems in Fig.2, while the spin splitting form in momentum space such as $p$-wave or $f$-wave is determined by the geometry of magnetic crystals. For example, as shown in Fig.2(a) and (b), there exists the symmetry $C_{2}||M_{yz}]$ with $M_{yz}$ being the mirror reflection perpendicular to $x$ axis as denoted by the black dashed line in the square lattice, as well as the symmetry $C_{2}||C_{6z}^{1}]$ in the hexagon lattice with the rotation $C_{6z}^{1}$ of $\tfrac{\pi}{3}$ radian around $z$ axis, implying a $p$-wave and $f$-wave spin splitting in momentum space, respectively.
In addition, we note that in the absence of nonmagnetic TRS then the symmetry $\bar{C}_{2}||T]$, the symmetry $C_{2}||\bar{E}]$ or $C_{2}||\bar{E}\bm{\tau}]$ is not always present with $\bar{E}$ or $\bar{E}\bm{\tau}$ connecting opposite-spin sublattices. For example, as shown in Fig.2(a), two sublattices in the square-lattice system have the same chirality, thus the uniform electric-magnetic field such as the light can not trigger opposite currents flowing on them, which is indispensable for the symmetry $C_{2}||\bar{E}\bm{\tau}]$. Instead, the opposite flux among the nearest-neighbor plaquettes is needed for realizing opposite sublattice currents. In contrast, for hexagon-lattice systems as shown in Fig.2(b), the reversed chirality between two sublattices enables opposite sublattice currents in the presence of the uniform electric-magnetic field, which explains why the odd-parity spin splitting has been observed in the hexagon-lattice AFM system under the light[33, 32].
@@
Bellow, the HH model is restudied to further reveal the odd-parity spin splitting in momentum space which reads
@@
Discussion and conclusion.—We first adopt the spin group method to identify sufficient conditions for the appearance of odd-parity ALM in collinear AFM’s. It is clarified that the breaking nonmagnetic TRS, enabling that the symmetry $\bar{C}_{2}||T]$ present in the even-parity ALM no longer holds, together with the symmetry $C_{2}||\bar{E}]$ or $C_{2}||\bar{E}\bm{\tau}]$ that connects opposite-spin sublattices trigger the odd-parity ALM characterized by the odd-parity spin splitting in real and momentum space. Furthermore, on account of the breaking nonmagnetic TRS induced by sublattice currents, HH model is systematically studied using the cluster slave spin method. Its phase diagram as a function of the Haldane hopping and onsite electron Coulomb repulsion is established, which consists of four regimes: AFMI at $\lambda=0$, CI with $C=2$, odd-parity ALM CI with $|M|>0$ and $C=2$, as well as the odd-parity ALMI with $|M|>0$ and $C=0$.

### 01 Apr 2026 v5

@@ Abstract
Following recent intensive studies on altermagnetism(ALM) characterized by non-relativistic even-parity spin splitting, realizing unconventional odd-parity magnetism has also attracted increasing interest. Here, using symmetry arguments based on spin-group analyses, we elucidate sufficient conditions for the emergence of odd-parity spin splitting in collinear antiferromagnetic systems, which is further established as the standard odd-parity ALM. It is derived that the odd-parity ALM arises from the following criteria: (i)the breaking nonmagnetic time reversal symmetry(TRS), i.e., the breaking real-space TRS; (ii)the long-range collinear compensated magnetism; (iii)the symmetry $C_{2}||\bar{E}]$ or $C_{2}||M]$ connecting opposite-spin sublattices, where $C_{2}$, $\bar{E}$, and $M$ respectively represent a $180^{\circ}$ rotation around the axis perpendicular to spins, the inversion, and the mirror reflection separating opposite-spin sublattices, directly reflecting the high-order harmonic($l\geq 3$) and the $p$-wave($l=1$) odd-parity ALM, respectively. Moreover, we utilize the well-known Haldane-Hubbard model to identify odd-parity spin splitting in the collinear ALM ground state, where (i)the nonmagnetic TRS is broken by opposite sublattice currents coming from the Haldane hopping; (ii)the symmetry $C_{2}||\bar{E}]$ is ensured because the currents flowing on opposite-spin sublattices are reversed.
@@
However, for collinear antiferromagnets with the breaking nonmagnetic TRS that can be caused by sublattice currents[38], light[33, 32], or the phase-dependent hopping coming from orbital orders[43], the symmetry $\bar{C}_{2}||T]$ appearing in even-parity ALM’s no longer holds. In contrast, the identity element in the nontrivial spin group $\bm{R}_{s}$ for collinear magnets, $\bar{C}_{2}||E]$, remains with $E$ being the identity in real space. Furthermore, as shown by Eq.(1), the odd-parity spin splitting in collinear magnets is ensured by the symmetry $C_{2}||C_{2z}]$, while for two-dimension systems, the operation $C_{2z}$ is equivalent to $\bar{E}$, such that the odd-parity spin splitting in two-dimension magnetic systems with long-range compensated collinear magnetism is ensured by the symmetry $C_{2}||\bar{E}]$. In addition, for $p$-wave altermagnets with spin splitting perpendicular to the mirror reflection $M$, which indicates that $C_{2}||M]E_{\bm{k}\sigma}=E_{M\bm{k}-\sigma}=E_{\bm{k}\sigma}$, the $p$-wave spin splitting is ensured by the symmetry $C_{2}||M]$. Therefore, in strong contrast to even-parity ALM’s, the long-range compensated collinear magnets exhibiting odd-parity spin splitting have the following characteristics: (i)the breaking nonmagnetic TRS then the breaking inversion symmetry of the Bravais lattice, i.e., $\bar{E}_{BL}$; (ii)the symmetry $C_{2}||\bar{E}]$ or $C_{2}||M]$, which directly enables that the nontrivial spin group expressed as $\bm{R}_{s}^{IV}=\{E||\bm{H}\}+\{C_{2}||A\bm{H}\}=\{E||\bm{H}\}+\{C_{2}||\bm{G}-\bm{H}\}$ has the distinct coset $A\bm{H}$ of the halved crystallographic Laue group $\bm{H}$ containing the inversion $\bar{E}$ or the mirror reflection $M$. To sum up, in spit of the similar form of the nontrivial spin group to even-parity ALM’s, the odd-parity spin splitting in compensated collinear magnets need to be classified into a new type, i.e., odd-parity altermagnetism, not only because the symmetry $\bar{C}_{2}||T]$ is absent, but also the inversion symmetry $\bar{E}$ or the mirror reflection $M$ is contained in the coset $A\bm{H}$.
@@
Haldane-Hubbard model.We use Haldane-Hubbard(HH) model[48] as an example where the nonmagnetic TRS is broken by sublattice currents to verify sufficient conditions for the occurrence of odd-parity ALM. As shown in Figs.2(a) and (b), the nonmagnetic TRS is broken by opposite currents flowing on two sublattices in the square- and hexagon-lattice magnetic system. In the square-lattice system[cf. Fig.2(a)], the reversed current flowing on opposite-spin sublattices enables the reversed chirality between them, then the compensated collinear magnetism triggers the emergence of a new form of symmetry $C_{2}||M_{yz}]$ with $M_{yz}$ being the mirror reflection perpendicular to $x$ axis as denoted by the black dashed line, which then gives rise to the $p$-wave ALM[38]. In addition, the reversed sublattice currents enables a new form of symmetry $C_{2}||\bar{E}_{BL}\bm{\tau}]$ with $\bm{\tau}=\hat{x},\hat{y}$ being the minimal vector connecting opposite-spin sublattices, indicating that this $p$-wave altermagnet has the time-reversal-symmetric energy band, i.e., $E_{\bm{k}\sigma}=E_{-\bm{k}-\sigma}$, in strong contrast to conventional $p$-wave magnets[31, 34, 35]. Therefore, the nontrivial spin group $\tilde{\bm{R}_{s}}$ in the square-lattice system can be expressed as $\tilde{\bm{R}}_{s}-[C_{2}||\bar{E}_{BL}\bm{\tau}]=[E||\bm{H}]+[C_{2}||M_{yz}\bm{H}]$ and $\bm{H}=\{E\}$. However, as shown in Fig.2(b), sublattice currents in the hexagon-lattice system have the same chirality between opposite-spin sublattices, thus the symmetry $C_{2}||C_{6z}]$ appears as the staggered magnetization sets in with $C_{6z}$ being the rotation of $\pi/3$ radian around the axis perpendicular to the plane. In addition, opposite-spin sublattices with the reversed current are connected by the symmetry $C_{2}||\bar{E}_{BL}\bm{\tau}]$ with $\bm{\tau}$ being the minimal vector connecting them. Therefore, the nontrivial spin group can be written as $\tilde{\bm{R}}_{s}-[C_{2}||\bar{E}_{BL}\bm{\tau}]=[E||\bm{H}]+[C_{2}||C_{6z}\bm{H}]$ and $\bm{H}=\{E,C_{6z}^{2},C_{6z}^{4}\}$, directly describing a $f$-wave spin splitting in real and momentum space.
To sum up, the nontrivial spin group for odd-parity ALM’s can be expressed as $\tilde{\bm{R}}_{s}=\bm{R}_{s}\oplus\{[\bar{C}_{2}||E],[C_{2}||\bar{E}_{BL}\bm{\tau}]\}$ with $\bar{E}_{BL}$ and $\bm{\tau}$ respectively being the inversion of the Bravais lattice and the minimal vector connecting opposite-spin sublattices, while the subgroup $\bm{R}_{s}$ reads $\bm{R}_{s}=\{E||\bm{H}\}+\{C_{2}||A\bm{H}\}=\{E||\bm{H}\}+\{C_{2}||\bm{G}-\bm{H}\}$ with $A$ being the inversion $\bar{E}$ or the mirror reflection $M$, respectively. Therefore, we have systematically studied the nontrivial spin group in the square- and hexagon-lattice HH model[cf. Fig2], where sufficient conditions for the odd-parity ALM are satisfied: (i)the nonmagnetic TRS is broken by opposite sublattice currents; (ii)the nontrivial spin subgroup reads $\bm{R}_{s}=\{E||\bm{H}\}+\{C_{2}||A\bm{H}\}$ with $A$ being the inversion $\bar{E}$ and the mirror reflection $M_{yz}$ in the hexagon- and square-lattice system, respectively.
@@
Bellow, HH model is restudied to further reveal the odd-parity spin splitting in momentum space which reads
@@
Discussion and conclusion.—We first adopt the spin group method to identify sufficient conditions for the appearance of odd-parity ALM in collinear AFM’s. It is clarified that the breaking nonmagnetic TRS, enabling that the symmetry $\bar{C}_{2}||T]$ present in the even-parity ALM no longer holds, together with the symmetry $C_{2}||\bar{E}]$ or $C_{2}||M]$ that connects opposite-spin sublattices trigger the odd-parity ALM characterized by the odd-parity spin splitting in real and momentum space. Furthermore, on account of the breaking nonmagnetic TRS induced by sublattice currents, HH model is systematically studied using the cluster slave spin method. Its phase diagram as a function of the Haldane hopping and onsite electron Coulomb repulsion is established, which consists of four regimes: AFMI at $\lambda=0$, CI with $C=2$, odd-parity ALM CI with $|M|>0$ and $C=2$, as well as the odd-parity ALMI with $|M|>0$ and $C=0$.